OpenStudy (anonymous):

you purchase a 50 meter roll of fencing, at a cost of $14.95 per meter, in order to build a dog run that is 340cm wide and 1380cm long. after you cut the four pieces of fencing from the roll, how much of fencing is left on the roll?

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) 1 meter = 100 centimeters So 50m x 100 = 5,000cm You have 5,000 centimeters 5,000 - 340 - 1,380 = 3,280cm 3,280cm is left on the roll OR 33.8 meters is left on the roll

OpenStudy (compassionate):

I would but you only accept messages from people who have fanned you (;

OpenStudy (dumbsearch2):

@Luis_Rivera It's LATEX.